OpenStudy (anonymous):

Simplify the following quotient of complex numbers into the form a + bi. -9+2i/-8-9i

OpenStudy (anonymous):

-9+2i/-8-9i =(-9+2i)(-8+9i) / (-8-9i)(-8+9i) <---multiply both numberator and denomintor with the conjugate of denominator so as to remove i in denomintator =(-9+2i)(-8+9i)/ (64+81) =(72-97i-18) / 145 =54/145 -(97/145)i

OpenStudy (anonymous):

thanks again. so u have to take the i from the denominator